find two numbers such that the sum of twice the first and the second is 92, and four times the first exceeds seven times the second by 2.
Answer by flame8855(424)

2x+3y=92
4x=7y+2
-4x-6y=-184
4x-7y=2
-13y=-182
y = 14
x= 25